please help me solve this question
Find a solution to the following system of equations.
-2x+5y=8
2x+2y=6

Answer by tasmart(9)

-2x+5y=8
2x+2y=6
7y=14 2x+2y=6
y=14/7 2x+2(2)=6
y=2 2x+4=6
2x=6-4
2x=2
x=1
check:-2(1)+5(2)=8
-2+10=8

2(1)+2(2)=6
2+4=6
y=2 x=1